OpenStudy (idealist10):

A bank pays interest continuously at the rate of 6%. How long does it take for a deposit of Q0 to grow in value to 2Q0?

OpenStudy (idealist10):

@ganeshie8 @SolomonZelman @campbell_st @eliassaab @mathmate @StudyGurl14 @tester97 @Ashleyisakitty @Jhannybean

OpenStudy (anonymous):

Find t so that \[ Q(t)= Q_0 \ e^{0.06 t}= 2 Q_0 \]

OpenStudy (anonymous):

This t does not depend on \( Q_0\)

OpenStudy (anonymous):

In fact you can show that \[ t=\frac {\ln 2}{.06}=11.5525 \]

OpenStudy (idealist10):

But the answer is \[\frac{ 50\ln 2 }{ 3 } yrs\]

OpenStudy (idealist10):

The answers are the same but is it okay if I leave the answer as ln(2)/0.06 instead of book's answer?

OpenStudy (anonymous):

They are the same.

OpenStudy (idealist10):

Thanks.